### v1

# Non-adaptive Bellman-Ford:

Abstract
The Bellman-Ford algorithm for single-source shortest paths repeatedly updates tentative distances in an operation called relaxing an edge. In several important applications a non-adaptive (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independent of the edge-weights. In a dense graph on $n$ vertices, the algorithm in its standard form performs $(1+o(1))n^{3}$ relaxations. An improvement by Yen from 1970 reduces the number of relaxations by a factor of two. We show that no further constant-factor improvements are possible, and every non-adaptive deterministic algorithm based on relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ steps. This improves an earlier lower bound of Eppstein of $(\frac{1}{6}-o(1))n^{3}$. Given that a non-adaptive randomized variant of Bellman-Ford with at most $(\frac{1}{3}+o(1))n^{3}$ relaxations (with high probability) is known, our result implies a strict separation between deterministic and randomized strategies, answering an open question of Eppstein.
1 Introduction
The Bellman-Ford algorithm [Bel58, For56, Moo59] is a classical method for computing single-source shortest paths in edge-weighted, directed graphs. Similarly to Dijkstra’s algorithm, the Bellman-Ford algorithm uses edge relaxations. Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distance from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$. Initially setting all tentative distances to +$\infty$+ ∞, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances. Relaxing an edge $(u,v)$ changes the tentative distance from $s$ to $v$ to its final (correct) value, if and only if the tentative distance from $s$ to $u$ is already correct and $(u,v)$ is the last edge of a shortest path from $s$ to $v$. Algorithms of this type are also known as label-correcting.
Besides correctly handling negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a non-adaptive (also called oblivious) way. This means that the sequence of edge-relaxations is fixed in advance, independent of the edge-weights of the input graph and independent of the outcomes of previous operations. This reduces overhead while attaining the same asymptotic worst-case bound, and makes the algorithm more amenable for implementation in parallel or distributed settings, e.g., in network routing protocols [Hed88, MR17].
It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$, if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in-between). Moreover, if for every possible path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed. Conversely, if the edges of some path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to 00 and all other weights set to $1$) the correct distance will not have been reached.
The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every possible path from $s$ to another vertex appears as a subsequence. The standard version of the Bellman-Ford algorithm relaxes edges in a round-robin fashion. In its non-adaptive implementation, this requires relaxing each edge $n-1$ times, where $n$ is the number of vertices. In a complete directed graph this leads to $(1+o(1))n^{3}$ relaxations and a proportional overall running time.
A classical improvement by Yen [Yen70] reduces the number of relaxations by a factor of two, while preserving the non-adaptive nature of the algorithm. Yen’s algorithm decomposes the edge-set of the input graph into forward edges and backward edges, both sets forming acyclic subgraphs. It then alternates between relaxing all forward- and all backward edges in topological order.
The question of whether a better relaxation order exists than the one given by Yen’s method has been open since 1970. Recently, Eppstein [Epp23] showed that every non-adaptive deterministic algorithm must do $(\frac{1}{6}+o(1))n^{3}$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement. Bannister and Eppstein [BE12] obtained a randomized improvement to Yen’s algorithm, with a high-probability guarantee on the algorithm’s success. In the non-adaptive setting, this improves Yen’s bound by a factor of $\frac{2}{3}$.
When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), better bounds are known (e.g., see [Epp23] and references therein), this setting is, however, not our concern in this paper. Lower bounds for adaptive algorithms have so far been elusive.
Our result.
We show that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ relaxation steps on a complete directed graph with $n$ vertices. This shows that Yen’s algorithm cannot be improved with the choice of a better relaxation sequence, by any constant factor. Moreover, since faster randomized approaches are known, randomization strictly helps for non-adaptive relaxation sequences. This answers an open question of Eppstein [Epp23].
2 Deterministic lower bound
Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V|=n$ and $|E|=n(n-1)$, and let $s\in V$ be a special vertex (the source). Let $w:E\rightarrow\mathbb{R}$ denote the edge-weights. Note that $G$ is simple, containing no loops or multiple edges.
The Bellman-Ford algorithm and its variants compute the distances from $s$ to every vertex in $V$ by initially setting $d(s,s)=0$ and $d(s,v)=+\infty$ for all $v\in V\setminus\{s\}$, and repeatedly performing operations $\operatorname{\mathsf{relax}}(e)$ for edges $e\in E$. The operation $\operatorname{\mathsf{relax}}(e)$ for $e=(u,v)$ sets $d(s,v)\leftarrow\min\{d(s,v),d(s,u)+w(e)\}$. The goal is to execute such steps until $d(s,v)$ is the correct distance from $s$ to $v$ for all $v\in V$. We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any $\operatorname{\mathsf{relax}}$ operation still reduces tentative distances after the algorithm has finished).
As we consider non-adaptive algorithms, the behaviour of the algorithm is fully specified by the sequence of relax operations $S=(e_{1},\dots,e_{L})$ it performs, where $e_{i}\in E$ for all $1\leq i\leq L$. We refer to $L=|S|$ as the length of the sequence.
All paths considered are simple. We say that a sequence $S$ of edges contains a path $(v_{1},\dots,v_{k})$ if $S$ contains $(v_{1},v_{2}),\dots,(v_{k-1},v_{k})$ as a subsequence (in the given order, but not necessarily contiguously). A $k$-path is a path $(s,v_{1},\dots,v_{k})$ for some set of $k$ distinct vertices $v_{1},\dots,v_{k}\in V\setminus\{s\}$. We say that a sequence of edges $S$ is $k$-valid, if it contains every possible $k$-path. By our earlier discussion, a sequence $S$ is a valid relaxation sequence in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.
Let $T_{k}$ denote the length of the shortest $k$-valid sequence for graph $G$.
Theorem 2.1.
$T_{k}\geq{(k-2)}{{n-1}\choose{2}}+(n-1)^{2}$, for all $12\leq k\leq n-$ 1.
Our claimed result follows by setting $k=n-1$ and observing $(n-3){{n-1}\choose{2}}+(n-1)^{2}\geq\frac{n^{3}}{2}-O(n^{2})$. In the remainder of the section we prove Theorem 2.1.
We proceed by induction on $k$. For $k=2$, the claim holds: to contain every $2$-path, a sequence must contain every edge except those pointing to $s$, and thus have length at least $(n-1)^{2}$.
Assume now $k>2$ and let $S=(e_{1},\dots,e_{T_{k}})$ be a shortest $k$-valid sequence. Observe that if $S$ contains every $k$-path, then it also contains every $(k-1)$-path, i.e., it is $(k-1)$-valid.
For each entry $e_{i}$ of $S$, call $e_{i}$ critical, if and only if some $h$-path $P$ with $h\leq k-1$ is contained in $(e_{1},\dots,e_{i})$ but not in $(e_{1},\dots,e_{i-1})$. In words, $e_{i}$ finishes $P$ and is the first edge in the sequence to do so. Otherwise, call $e_{i}$ non-critical.
Let $N$ be the number of non-critical entries in $S$. Observe that $T_{k}=|S|\geq T_{k-1}+N$. This is because, as the non-critical edges are not needed for finishing any prefix of a $(k-1)$-path, the subsequence of $S$ obtained by removing all non-critical edges is still $(k-1)$-valid, and thus of length at least $T_{k-1}$.
Let $q_{i}$ be the endpoint of $e_{i}$, i.e., $e_{i}=(x,q_{i})$, for some $x,q_{i}\in V$. Let $Q_{i}$ denote the set of endpoints of critical edges in the sequence $(e_{i},\dots,e_{T_{k}})$. More precisely, set $Q_{T_{k}+1}=\emptyset$, and for $i=T_{k},\dots,1$, set $Q_{i}=Q_{i+1}\cup\{q_{i}\}$, if $e_{i}$ is critical, and set $Q_{i}=Q_{i+1}$ if $e_{i}$ is non-critical. Let us also denote, $\bar{Q}_{i}=(V\setminus\{s\})\setminus Q_{i}$.
Call a critical entry $e_{i}$ supercritical whenever $|Q_{i}|>|Q_{i+1}|$. In words, such an edge $e_{i}$ adds the new endpoint $q_{i}$ to the set of critical endpoints as we go right to left. As $S$ contains paths ending in every vertex (other than the source), notice that $Q_{1}=V\setminus\{s\}$. This means that there are $n-1$ supercritical entries in $S$. Denote these (left-to-right) by $e_{i_{1}},\dots,e_{i_{n-1}}$, with their (distinct) endpoints $q_{i_{1}},\dots,q_{i_{n-1}}$.
[…]
Repeating the argument for each supercritical edge, we conclude that all edges $(q_{i_{b}},q_{i_{a}})$ for $11\leq a<b\leq n-$ 1 must appear as non-critical entries. This implies a lower bound on the number of non-critical entries $N\geq{{n-1}\choose{2}}$, from which $T_{k}=|S|\geq T_{k-1}+{{n-1}\choose{2}}$ follows. Using the induction hypothesis to lower bound $T_{k-1}$ yields the claimed result.