### v2

# Non-adaptive Bellman-Ford:

Abstract
The Bellman-Ford algorithm for single-source shortest paths repeatedly updates tentative distances in an operation called relaxing an edge. In several important applications a non-adaptive (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independently of the edge-weights. Such an implementation performs, in a dense graph on $n$ vertices, $(1+o(1))n^{3}$ relaxations. An improvement by Yen from 1970 reduces the number of relaxations by a factor of two. We show that no further constant-factor improvements are possible, and every non-adaptive deterministic algorithm based on relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ steps. This improves an earlier lower bound of Eppstein of $(\frac{1}{6}-o(1))n^{3}$. Given that a non-adaptive randomized variant of Bellman-Ford with at most $(\frac{1}{3}+o(1))n^{3}$ relaxations (with high probability) is known, our result implies a strict separation between deterministic and randomized strategies, answering an open question of Eppstein. On the complexity side, we show that deciding whether a given relaxation sequence is guaranteed to yield correct distances is NP-hard, even with the complete graph as input.
1 Introduction
The Bellman-Ford algorithm [Bel58, For56, Moo59] is a classical method for computing single-source shortest paths in edge-weighted, directed graphs. Similarly to Dijkstra’s algorithm, the Bellman-Ford algorithm uses edge relaxations. Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distances from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$. Initially setting the tentative distance from $s$ to $s$ itself to 00 and to all other vertices to +$\infty$+ ∞, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances. Relaxing an edge $(u,v)$ changes the tentative distance from $s$ to $v$ to its final (correct) value, if and only if the tentative distance from $s$ to $u$ is already correct and $(u,v)$ is the last edge of a shortest path from $s$ to $v$. Algorithms of this type are also known as label-correcting.
Besides the correct handling of negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a non-adaptive (also called oblivious) way. This means that the sequence of edge-relaxations is fixed in advance, independently of the edge-weights of the input graph and of the outcomes of previous operations. This reduces overhead while maintaining the same asymptotic worst-case running time. A non-adaptive sequence of edge-relaxations also makes the algorithm amenable for implementation in parallel or distributed settings, e.g., in network routing protocols [Hed88, MR17], where centralized control may not be available.
It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$ if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in between). If for every simple path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed. Conversely, if the edges of some simple $s$-to-$v$ path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to 00 and all other weights set to $1$) the correct distance will not have been reached.
The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every simple path from $s$ to another vertex appears as a subsequence. The standard non-adaptive version of the Bellman-Ford algorithm relaxes all $m$ edges in a round-robin fashion $n-1$ times, where $n$ is the number of vertices. The resulting sequence of edges clearly contains every simple path, and in a complete directed graph the number of edge-relaxations is $(1+o(1))n^{3}$, with a proportional overall running time.
A classical improvement by Yen [Yen70] reduces the number of relaxations by a factor of two, while preserving the non-adaptive nature of the algorithm. Yen’s algorithm decomposes the edge set of the input graph into forward edges and backward edges, both sets forming acyclic subgraphs. It then alternates between relaxing all forward- and all backward edges in topological order.
The question of whether a better relaxation order exists than the one given by Yen’s method has been open since 1970. Recently, Eppstein [Epp23] showed that every non-adaptive deterministic algorithm must perform $(\frac{1}{6}-o(1))n^{3}$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement. Bannister and Eppstein [BE12] obtained a randomized improvement to Yen’s algorithm, with a high-probability guarantee on the algorithm’s success. In the non-adaptive setting, this improves Yen’s bound by a factor of $\frac{2}{3}$.
Our results.
We show (§ 2) that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ relaxation steps on a complete directed graph with $n$ vertices. This shows that Yen’s algorithm cannot be improved with the choice of a better relaxation sequence, by any constant factor. Moreover, since faster randomized approaches are known, randomization strictly helps for non-adaptive relaxation sequences. This answers an open question of Eppstein [Epp23].
On the complexity side, two natural questions arise. Given a directed graph $G$ and a source vertex $s$, we may ask for a shortest non-adaptive sequence of edge-relaxations that compute the correct distances from $s$, for all possible configurations of edge-weights. In light of the previous discussion, this amounts to finding a shortest sequence of edges that contains as subsequence every simple path from $s$ in $G$. A second question is to decide, whether a given sequence $S$ of edges is a valid relaxation sequence, i.e., whether it contains every simple path from $s$ in $G$.
While we suspect both problems to be (NP-)hard, we can only show this for the second, and leave the first one open. The hardness of the second problem holds even for the complete graph $G$ (§ 3).
We note that both problems admit polynomial-time algorithms in the easier case of BF-orderable graphs, i.e., graphs in which every edge needs to be relaxed at most once; this results from 1980s work of Mehlhorn and Schmidt [MS86], and Haddad and Schäffer [HS88].
Further related work.
Due to the fundamental nature of the single-source shortest paths problem in the presence of negative weights, several attempts were made to establish its precise complexity or to improve the long-standing $O(mn)$ runtime of the Bellman-Ford algorithm. We refer to Eppstein [Epp23] for a discussion of results in different models and for variants of the problem. Note that complexity bounds are sensitive to the exact modeling assumptions, in particular, they differ between the non-adaptive and adaptive settings.
When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), a recent breakthrough by Fineman [Fin24] has reduced the complexity to $\tilde{O}(mn^{8/9})$, with a later improvement [HJQ25] to $\tilde{O}(mn^{4/5})$. These results may explain why lower bounds for adaptive algorithms have been elusive. Nevertheless, an asymptotic lower bound of $\Omega(n^{3})$ has recently been shown [AHS+24], for adaptive relaxation-based algorithms with a certain restricted set of operations (while classical Bellman-Ford-variants fall within these restrictions, the mentioned recent improvements do not).
2 Deterministic lower bound
Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V|=n$ and $|E|=n(n-1)$, and let $s\in V$ be the source vertex. Let $w:E\rightarrow\mathbb{R}$ denote the edge-weights. Note that $G$ is simple, containing no loops or multiple edges.
The Bellman-Ford algorithm and its variants compute the distances from $s$ to every vertex in $V$ by initially setting $d(s,s)=0$ and $d(s,v)=+\infty$ for all $v\in V\setminus\{s\}$, and repeatedly performing operations $\operatorname{\mathsf{relax}}(e)$ for edges $e\in E$. The operation $\operatorname{\mathsf{relax}}(e)$ for $e=(u,v)$ sets $d(s,v)\leftarrow\min\{d(s,v),d(s,u)+w(e)\}$. The goal is to execute such steps until $d(s,v)$ is the correct distance from $s$ to $v$ for all $v\in V$. We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any relax operation still reduces tentative distances after the algorithm has finished).
As we consider non-adaptive algorithms, the behaviour of the algorithm is fully specified by the sequence of relax operations $S=(e_{1},\dots,e_{L})$ it performs, where $e_{i}\in E$ for all $1\leq i\leq L$. We refer to $L=|S|$ as the length of the sequence. Our main result is the following.
Theorem 2.1.
Every valid relaxation sequence for $G$ must have length at least $(\frac{1}{2}-o(1))n^{3}$.
In the remainder of the section we prove the theorem. All paths considered are simple. We say that a sequence $S$ of edges contains a path $(v_{1},\dots,v_{k})$ if $S$ contains $(v_{1},v_{2}),(v_{2},v_{3}),\dots,(v_{k-1},v_{k})$ as a subsequence (in the given order, but not necessarily contiguously). A $k$-path is a path $(s,v_{1},\dots,v_{k})$ for some set of $k$ vertices $\{v_{1},\dots,v_{k}\}\subseteq V\setminus\{s\}$. We say that a sequence of edges $S$ is $k$-valid, if it contains every possible $k$-path. By our earlier discussion, a sequence $S$ is a valid relaxation sequence for $G$ in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.
The task of finding a short $k$-valid sequence appears related to the task of finding a short permutation supersequence, a well-studied question in combinatorics. In that context, $c_{n}$ denotes the smallest possible length of a sequence over $\{1,\dots,n\}$ that contains each permutation of $\{1,\dots,n\}$ as a subsequence (not necessarily contiguously).
It is known that $c_{n}\leq\lceil n^{2}-\frac{7}{3}n+\frac{19}{3}\rceil$ [Rad12], see also the survey of Engen and Vatter [EV20]. Most relevantly for us, a lower bound of the form $c_{n}\geq(1-o(1))n^{2}$ was shown in 1976 by Kleitman and Kwiatkowski.
Lemma 2.2 ([KK76]).
For all $n\geq 1$ and all $\varepsilon>0$, we have $c_{n}>n^{2}-C_{\varepsilon}\cdot n^{7/4+\varepsilon}$, with $C_{\varepsilon}$ depending only on $\varepsilon$.
We will use Lemma 2.2 to prove Theorem 2.1.111In an earlier version of the paper we gave a more direct proof of Theorem 2.1, but the proof was faulty.
It is well-known that whenever $N$ is even, the edge-set of a complete (undirected) graph on $N$ vertices can be decomposed into $N-1$ disjoint perfect matchings, i.e., sets of $\frac{N}{2}$ non-incident edges. (Such a decomposition is called a $1$-factorization, often studied in the context of round-robin tournaments, e.g., see [MR85]). If $N$ is odd, we can simply ignore one of the vertices and decompose the remainder of the graph into $N-2$ disjoint matchings, each consisting of $\frac{N-1}{2}$ edges.
Denote the vertex set of the complete directed graph $G$ as $V=\{s,v_{1},\dots,v_{n-1}\}$. Consider a complete (undirected) graph on $\{v_{1},\dots,v_{n-1}\}$. By the previous paragraph, regardless of the parity of $n$, we can find in this graph $n-3$ disjoint matchings, each of size $\frac{n-2}{2}$. Denote such a collection of matchings as $E_{1},\dots,E_{n-3}$ (each $E_{i}$ is a set of undirected edges).
For each matching $E_{i}$ we form two sets of directed edges in $G$; one where each edge of $E_{i}$ is oriented from smaller to larger index vertex (call this $F_{i}$), and one where each edge of $E_{i}$ is oriented from larger to smaller index vertex (call this ${H_{i}}$). We thus obtain a collection =$\mathcal{F}=\{F_{1},H_{1},\dots,F_{n-3},{H_{n-3}}\}$ of pairwise disjoint sets, each consisting of pairwise non-incident directed edges in $G$.
Let $S$ be a shortest $(n-1)$-valid, or simply valid, sequence of edges for graph $G$. For any $F\in\mathcal{F}$ let $S_{F}$ denote the maximal subsequence of $S$ consisting of elements in $F$.
We claim that $S_{F}$ must contain every permutation of $F$ as a subsequence, and thus the lower bound of Lemma 2.2 applies to $S_{F}$.
Indeed, suppose some permutation of $F$ is not contained in $S_{F}$, and denote this permutation (of directed edges of $G$) as $(e_{1},\dots,e_{\frac{n-2}{2}})$, where $e_{i}$ is the directed edge $(x_{i},y_{i})$, for $x_{i},y_{i}\in V\setminus\{s\}$. In this case, $S$ could not contain the sequence of edges $((s,x_{1}),\mathbf{e_{1}},(y_{1},x_{2}),\mathbf{e_{2}},(y_{2},x_{3}),\dots,\mathbf{e_{\frac{n-2}{2}}})$ that form a simple path from the source to $y_{\frac{n-2}{2}}$, and thus $S$ could not be valid.
Since $|F|=\frac{n-2}{2}$ for all $F\in\mathcal{F}$, we have $|S_{F}|\geq(\frac{1}{4}-o(1))n^{2}$, by Lemma 2.2. As edge sets in $\mathcal{F}$ are pairwise disjoint, the corresponding subsequences of $S$ do not overlap. Thus, we have
$|S|~{}\geq~{}|\mathcal{F}|\cdot\left(\frac{1}{4}-o(1)\right)n^{2}~{}\geq~{}\left(2n-6\right)\left(\frac{1}{4}-o(1)\right)n^{2}~{}\geq~{}\left(\frac{1}{2}-o(1)\right)n^{3}.$
This completes the proof. $\hfill\qed$
3 Hardness of deciding validity
Again, let $G$ be a directed graph, with a special vertex $s$. Let $S=(e_{1},\dots,e_{|S|})$ be a sequence of edges of $G$. The relaxation validity question asks whether $S$ contains as subsequence every simple path of $G$ starting in $s$, in other words, whether $S$ is a valid relaxation sequence.
Theorem 3.1.
Relaxation validity is co-NP-complete.
In the remainder of the section we prove Theorem 3.1. Membership in co-NP is trivial; if $S$ is not valid, then some simple path $P$ witnesses this. Verifying that $S$ does not contain $P$ can be done by greedily matching edges of $P$ in their order from $s$, to the earliest occurrence in $S$, to the right of the entry matched to the predecessor edge.
To show co-NP-hardness, we reduce from the all-permutations supersequence (APS) problem, shown to be co-NP-hard by Uznański [Uzn15]. APS asks whether a given sequence $T=(t_{1},\dots,t_{|T|})$ contains every permutation of $\{1,\dots,n\}$ as a subsequence. (Alternatively, one may ask whether there is a permutation not contained in $T$, which is consequently NP-hard).
The reduction from APS to relaxation validity is straightforward. Consider a complete directed graph $G$ with a set of $12n+$ 1 vertices denoted $V=\{s,x_{1},y_{1},\dots,x_{n},y_{n}\}$. For $i\in\{1,\dots,n\}$, let $e_{i}$ denote the special edge $(x_{i},y_{i})$.
Let $Q$ be a valid relaxation sequence of $G$, i.e., a $2n$-valid sequence of edges. Let $R$ be the sequence of edges obtained from $Q$ by deleting every occurrence of $e_{i}$ for all $i$.
We construct a relaxation sequence $S$ of $G$ as follows. For $i=1,\dots,|T|$, we form $S_{i}$ by appending to the end of sequence $R$ the edge $e_{j}$, for $j=t_{i}$. Then, we concatenate $S_{1},\dots,S_{|T|}$, and finally $R$, to form the sequence $S$. From the preceding discussion, we can enforce $|Q|,|R|\in O(n^{3})$ and the construction of $S$ takes time polynomial in $|T|$. (Notice that we do not require $Q$ to be optimal, a simple round-robin sequence of length $O(n^{3})$ is sufficient).
We claim that $T$ contains every permutation of $\{1,\dots,n\}$ if and only if $S$ contains every simple path starting from $s$ in $G$, from which Theorem 3.1 follows.
First, suppose $T$ does not contain some permutation $\pi_{1},\dots,\pi_{n}$. Then, $S$ does not contain the sequence of edges $(e_{\pi_{1}},\dots,e_{\pi_{n}})$ as these edges only appear in $S$ in correspondence to entries of $T$. Consequently, $S$ cannot correctly relax the path $(s,x_{\pi_{1}},y_{\pi_{1}},\dots,x_{\pi_{n}},y_{\pi_{n}})$.
Suppose now that $T$ contains every permutation of $\{1,\dots,n\}$. Then, every simple $k$-path $P$ is correctly relaxed by $S$. Indeed, since $T$ contains every permutation, $S$ contains the (possibly incomplete) permutation of edges $e_{\pi_{1}},e_{\pi_{2}},\dots$ in $P$, in their order of appearance in $P$. The portions of $P$ between any two edges $e_{\pi_{i}}$, $e_{\pi_{i+1}}$ or before the first, respectively, after the last such edge, are necessarily contained in some copy of $R$ inserted between the corresponding occurrences of $e_{\pi_{i}}$, $e_{\pi_{i+1}}$ in $S$, resp., at the beginning or at the end of $S$. This concludes the proof.
4 Conclusions
We showed that the number of edge-relaxations in Yen’s variant of the Bellman-Ford algorithm cannot be improved in a deterministic non-adaptive way by any constant factor. Our lower bound is shown for the complete directed graph as the input. Besides serving as a worst-case example, the complete graph is also natural as it captures full non-adaptivity: If an algorithm is oblivious to both the weights and the structure of the graph, then it must attempt to relax every possible edge, as if the input were a complete graph.
In the randomized case, a gap between the $(\frac{1}{12}-o(1))n^{3}$ lower bound due to Eppstein and the $(\frac{1}{3}+o(1))n^{3}$ upper bound due to Bannister and Eppstein remains [Epp23, BE12]. For incomplete dense graphs, i.e., with the number of edges $m\geq n^{1+{\varepsilon}}$, Eppstein shows that in the worst case, the $O(mn)$ bound cannot be improved by non-adaptive relaxation sequences (in this case, the algorithm knows the graph structure but not the weights). For sparse families, a gap between the $O(mn)$ upper bound and an $\Omega(mn/\log{n})$ lower bound remains [Epp23].
The complexity of finding a short relaxation sequence for a given graph remains open. The problem appears similar to the shortest supersequence problem, known to be NP-hard in several variants (e.g., see [Mai78, RU81]). The difficulty of extending these results to relaxation sequences is twofold. The paths in our setting are simple, i.e., without repeated vertices or edges, whereas in the inputs to the shortest supersequence problem symbols can appear multiple times. To our knowledge, the complexity of finding the shortest supersequence of a set of (partial) permutations is open. A second difficulty is that the set of input paths in our setting is only implicitly given and the number of paths is, in general, exponential.